OpenStudy (anonymous):

Evaluate the expression when p = –4 and q = 2. 20 + 6q ÷ p A. –8 B. 17 C. –17 D. 8

OpenStudy (anonymous):

deez nuts

OpenStudy (anonymous):

substitude the values of x and y in the expression \[20+6*(2)/(-4)\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

do the multiplication first

OpenStudy (anonymous):

6*2=12 then do the division

OpenStudy (texaschic101):

its just a matter of substitution...and then PEMDAS 20 + 6(2) / -4 multiplication, then division, then addition

OpenStudy (anonymous):

12/(-4)=-3 now do the addition

OpenStudy (anonymous):

20+(-3)=20-3=17

OpenStudy (anonymous):

6*2=12 20+12/-4 12/-4=-3 20+-3=17 answer is b